What is the pH of water with an inert electrolyte
Answer #1

pH = -log(aH)

where aH is activity of H+

aH = [H+] γH+

Activity coefficient, γH+ = 0.8

For water at 25 °C, [H+] = 10-7 M

aH = 0.8 * 10-7 M

pH = - log(0.8 * 10-7) = 7.097